# Why Reasoning Matters? A Survey of Advancements in Multimodal Reasoning (v1)

### v1

@@ 1 Introduction
Reasoning is a fundamental aspect of human intelligence, enabling us to solve complex problems effectively. In LLMs, reasoning has shown promising capabilities, leading to significant advancements in various domains (Bi et al., 2024b), including arithmetic, commonsense, and symbolic reasoning. For instance, Chain-of-Thought (CoT) prompting mimics human stepwise problem-solving to boost LLM performance, especially in very large models.
Beyond CoT, other prompting strategies, such as generated knowledge prompting and tree search algorithms (e.g., STAR search (Zelikman et al., 2022)), have been incorporated into LLM training to further enhance reasoning capabilities. These methods encourage the model to articulate its internal reasoning in natural language, which reinforces its understanding of the task. By laying out intermediate steps, the model can verify and adjust its logic before arriving at the final answer.
@@
Initial research has demonstrated that integrating Chain-of-Thought (CoT) reasoning across different modalities can significantly enhance the performance of MLLMs. For example, Yao et al. (2024); Zhang et al. (2024a); Bi et al. (2024a) reports that the LLama series models’ reasoning abilities are greatly improved by solely utilizing filtered self-generated reasoning paths.
@@ 3 Post-training improvements
$s_{t}$ is the sequence of tokens up to time $t$, and $a_{t}$ is the next token. The advantage function $A(s_{t},a_{t})$ estimates the benefit of choosing $a_{t}$ in state $s_{t}$, guiding the model toward higher-utility token sequences. This objective can be directly optimized with respect to the model parameters using policy gradient methods (Williams, 1992). Notably, this formulation enables the training of LLMs not merely to predict the next token, but to generate complete reasoning trajectories that are optimized for long-term rewards.
[…]
@@ 3.1 Policy Optimization
To solve the above reward maximization problem, recent advancements in visual reasoning have prominently focused on integrating reinforcement learning (RL) and imitation learning (IL) to align MLLMs more closely with human reasoning. IL approaches, such as Thought Cloning (Wei et al., 2025), surpass traditional cloning by aligning intermediate reasoning steps instead of final outputs alone. Coupled with iterative methods like DAgger, this reduces dataset shift and hallucination. LLaVA-Critic (Xiong et al., 2024) further generalizes reward signals via preference-based alignment, boosting multimodal reasoning effectiveness with minimal modality-specific fine-tuning. RL from Simulations (RLS3) (Waite et al., 2025), employing Soft Actor-Critic (SAC) agents, significantly enhances spatial reasoning accuracy and efficiency in models like PaliGemma, outperforming random exploration baselines (Waite et al., 2025). FuRL (Fu et al., 2024c) addresses RL reward misalignment through dual-stage alignment, effectively mitigating sparse reward challenges. Adaptive Reward Programming (ARP-DT+) (Kim et al., 2023), integrating CLIP representations with Decision Transformers, enhances reasoning robustness across domains by employing active inference to counter distractions and domain shifts.
@@ 3.2 Reward Alignment
Recent approaches have explored iterative refinement and reflection mechanisms, token-level rewards, and automated benchmarking to enhance reasoning quality and mitigate hallucinations. PARM++ (Guo et al., 2025) introduces a reflection-based mechanism that iteratively refines outputs by identifying misalignments and actively correcting them through self-assessment loops, significantly improving test-time GenEval scores up to three reflection cycles (Guo et al., 2025). Similarly, REVERIE employs rationale-based training, explicitly supervising reasoning steps, thus improving coherence and reducing hallucinations compared to models without rationales (Zhang et al., 2024f). Fine-grained reward models, such as MMViG (Yan et al., 2024) and Reward Alignment via Preference Learning (RAPL) (Tian et al., 2024), emphasize granular feedback at individual reasoning steps, achieving precise improvements without manual reward tuning. Similarly, CLIP-DPO (Ouali et al., 2024) leverages pre-trained CLIP directly as a reward function, streamlining reward modeling. Challenges such as constrained perceptual fields are addressed through dynamic evaluation strategies like zooming and shifting, emphasizing the importance of sequential decision-making in visual reasoning (Wang et al., 2024j). Structural issues like positional bias and patch-boundary effects further underscore the need for targeted architectural optimizations (Zhang et al., 2024d). Recent approaches integrate reward signals directly into generation. TLDR (Fu et al., 2024a) applies token-level binary rewards, enabling real-time feedback to mitigate hallucinations and boost both annotation efficiency and backbone model performance (Fu et al., 2024a). Complementarily, EACO enhances intermediate object recognition, significantly reducing hallucinations compared to earlier models like LLaVA-v1.6-7B (Wang et al., 2024g). Distillation frameworks also contribute: FIRE leverages iterative student–teacher feedback and structured evaluation to refine visual reasoning (Li et al., 2024d), while SILKIE demonstrates the scalability of GPT-4V-generated feedback in improving perception and cognition without human intervention (Li et al., 2023b).
@@ 3.3 Model Architecture
Nested architectures, such as MaGNeTS, have emerged as effective means of balancing computational efficiency with model accuracy by employing parameter-sharing and caching mechanisms during inference (Goyal et al., 2025). Intrinsic activation approaches, notably ROSS, bypass external modules by integrating multimodal understanding directly into the model’s core, yielding improved adaptability to diverse visual inputs, including depth maps (Wang et al., 2024c). Additionally, DIFFLMM incorporates diffusion models and attention-based segmentation, enhancing visual grounding precision (Cao et al., 2024a), while Mini-Monkey targets the ”sawtooth effect” in lightweight models through adaptive cropping and scale compression, achieving superior visual-text alignment without extensive computational resources (Huang et al., 2024b). SEA dynamically adjusts embedding alignment strategies to maintain performance across varying resolutions and model sizes (Yin et al., 2024). Concurrently, PAE-LaMM demonstrates the combined effectiveness of vision encoder adaptation and pixel reconstruction tasks, systematically enhancing visual detail awareness and question-answering performance (Gou et al., 2024). Conversely, LLAVIDAL and EventLens uniquely integrate domain-specific visual elements, such as object-centric embeddings and event-aware tokens, respectively, directly within LLM architectures to effectively reason about activities and temporal contexts (Reilly et al., 2024; Ma et al., 2024a).
@@ 3.4 Spatial Temporal Modeling
Unlike pure language reasoning, visual reasoning involves a richer action space, allowing operations such as zooming, cropping, and frame selection, which enables the model to interact with and manipulate visual inputs. The GeoGLIP pipeline utilizes geometric pre-training with a dynamic feature router to advance fine-grained visual mathematical understanding by balancing symbolic and visual modalities based on task demands (Zhang et al., 2025). Concurrently, hallucination issues are mitigated via targeted fine-tuning and prompting strategies grounded in recent vision-language work like CapQA (Hu et al., 2025). For spatial reasoning, simpler graph topologies have been shown to enhance performance, while structural complexity reduces accuracy in edge-centric tasks (Zhu et al., 2024). Augmentation via uncertainty-aware active inference further boosts attribute detection precision (Zhang et al., 2024e). Intrinsic spatial-temporal modeling is advanced through architectures embedding multi-modal understanding without reliance on external depth tools (Wang et al., 2024c), and dual-branch structures improve video temporal grounding in multi-hop reasoning (Chen et al., 2024f). The shift toward multimodal benchmarks better reflects real-world spatial-temporal reasoning (Li et al., 2024a). Integrating rationalization with answer generation enhances temporal reasoning (Zohar et al., 2024), while automated annotation improves data scalability and bridges gaps between commercial and open-source Video-LLMs (Li et al., 2024e). Multi-modal integration and active inference optimize modality selection for multi-step spatial-temporal reasoning (Zhang et al., 2024l). Adjusting learning objectives to minimize hallucinations improves precision and EOS decision-making (Yue et al., 2024).
@@ 4 Test-Time compute
[…]
@@ 4.1 Search Strategies
Monte Carlo Tree Search (MCTS) has emerged as a powerful tool for managing uncertainty and sequential decision-making. Imam et al. (2025) applied MCTS to enhance temporal reasoning, reducing hallucinations in time-dependent data. Building on this, CoMCTS(Yao et al., 2024) unified multiple MLLMs within a collaborative tree search, addressing model bias and improving accuracy and efficiency. Vision-specific strategies advanced with VisVM (Wang et al., 2024f), which estimates long-term candidate value to reduce hallucinations, outperforming immediate-alignment methods like CLIP-PRM. Video reasoning similarly benefited from reward modeling and adaptive exploration to prioritize key frames (Yang et al., 2024b). Structured search methods also gained traction. LLaVA-o1 (Xu et al., 2025) uses stage-level beam search for complex reasoning, while MVP (Qu et al., 2024) aggregates certainty across multi-perspective captions to resist adversarial inputs. DC2 (Wang et al., 2024e) applies MCTS-based cropping to focus on salient image regions for high-resolution reasoning. Multimodal and temporal search frameworks like VideoVista (Li et al., 2024e), WorldRetriver (Zhang et al., 2024l), and DynRefer (Zhao et al., 2024) surpass static baselines using adaptive sampling, fusion, and stochastic inference. Step-by-step comparative reasoning with LMs also enhances video QA (Nagarajan & Torresani, 2024). Innovations in context refinement (VURF (Mahmood et al., 2024)), image decomposition (V*(Wu & Xie, 2023)), and dynamic tool use (AVIS(Hu et al., 2023)) highlight the shift toward adaptive visual reasoning. FuRL (Fu et al., 2024c) aligns reward modeling with iterative fine-tuning for better performance. In embodied AI, combining chain-of-thought, self-verification, and MCTS-driven planning enables scalable, robust decision-making in dynamic environments (Shin et al., 2024).
@@ 4.2 Adaptive Inference
Adaptive inference is reshaping vision-language reasoning by enabling dynamic, context-sensitive processing that improves both accuracy and efficiency. Central to this is the iterative evaluation and refinement of outputs, often through internal feedback and external verification. LOOKBACK enhances correction accuracy through iterative visual re-examination within each reasoning step (Wu et al., 2024b). IXC-2.5 balances performance and response length via Best-of-N sampling guided by reward models (Zang et al., 2025), while PARM++ uses reflection-based refinement to align generated images with prompts (Guo et al., 2025). Further innovations include ProgressCaptioner’s sliding-window approach for tracking action progress over time (Xue et al., 2024). To combat multi-modal hallucinations, MEMVR triggers visual retracing based on uncertainty, optimizing cost-accuracy tradeoffs (Zou et al., 2024a), while MVP aggregates certainty across diverse views (Qu et al., 2024). SID applies token-level contrastive decoding to filter irrelevant content (Huo et al., 2024). Models like DualFocus integrate macro and micro reasoning for fine-grained attention control (Cao et al., 2024b), and LISA++ enables test-time compute scaling without retraining (Yang et al., 2024a). PerceptionGPT accelerates inference via adaptive token embeddings that encode spatial cues and dynamically weigh layers (Pi et al., 2023).
@@ 4.3 Reward
Guo et al. (2025) introduce ORM and PRM, showing that ORM’s final-step evaluation significantly enhances image generation through ranking data. PARM++ further refines this method by enabling iterative reflection for self-correction, thus improving prompt fidelity. Complementarily, Hao et al. (2025) highlight persistent errors in VLM spatial reasoning and explanation, proposing reward re-ranking as an interim solution. To assess VL-GenRMs more comprehensively, Li et al. (2024b) propose VL-RewardBench, pinpointing perception and reasoning challenges while advocating critic training and co-evolutionary reward learning. Self-training methods also gain prominence: R3V (Cheng et al., 2024b) leverages synthesized rationales to enhance noise-robust Chain-of-Thought (CoT) refinement, while CECE (Cascante-Bonilla et al., 2024) enriches evaluation using LLM-generated entailment/contradiction captions. Enhanced supervision methods also emerge. Zhang et al. (2024k) enhance CoT reasoning through ShareGPT-4O distillation, Supervised Fine-Tuning (SFT), and Direct Preference Optimization (DPO). LLaVA-Critic (Xiong et al., 2024) builds upon this using critic-guided DPO without external human feedback. FuRL (Fu et al., 2024c) introduces a two-stage tuning approach addressing misalignments and temporal errors, while HYDRA (Ke et al., 2024) employs dynamic loops for adaptive instruction ranking. Additionally, m&m’s (Ma et al., 2024b) provides a structured framework for evaluating multi-step planning via tool and argument accuracy.

### v2

@@ 1 Introduction
Reasoning is a fundamental aspect of human intelligence, enabling us to solve complex problems effectively. In LLMs, reasoning has shown promising capabilities, leading to significant advancements in various domains (Bi_2024), including arithmetic, commonsense, and symbolic reasoning. For instance, Chain-of-Thought (CoT) prompting mimics human stepwise problem-solving to boost LLM performance, especially in very large models.
Beyond CoT, other prompting strategies, such as generated knowledge prompting and tree search algorithms (e.g., STAR search (2203.14465-zelikman_star:_2022)), have been incorporated into LLM training to further enhance reasoning capabilities. These methods encourage the model to articulate its internal reasoning in natural language, which reinforces its understanding of the task. By laying out intermediate steps, the model can verify and adjust its logic before arriving at the final answer.
@@
Initial research has demonstrated that integrating Chain-of-Thought (CoT) reasoning across different modalities can significantly enhance the performance of MLLMs. For example, 2412.18319-yao_mulberry:_2024; 2410.02884-zhang_llama-berry:_2024; bi2024unveilingvisualperceptionlanguage reports that the LLama series models’ reasoning abilities are greatly improved by solely utilizing filtered self-generated reasoning paths.
@@ 3 Post-training improvements
$s_{t}$ is the sequence of tokens up to time $t$, and $a_{t}$ is the next token. The advantage function $A(s_{t},a_{t})$ estimates the benefit of choosing $a_{t}$ in state $s_{t}$, guiding the model toward higher-utility token sequences. This objective can be directly optimized with respect to the model parameters using policy gradient methods (williams1992simple). Notably, this formulation enables the training of LLMs not merely to predict the next token, but to generate complete reasoning trajectories that are optimized for long-term rewards.
ȷ Policy optimization is still in its early stages, often relying on existing methods with a primary focus on text token optimization. However, approaches from visual reinforcement learning are worth exploring to better align policies with perceptual and spatial dynamics. Reward alignment remains heavily reliant on textual signals, highlighting an opportunity to shift focus toward visual-based rewards. Developing methods that can interpret and adapt to visual feedback—such as spatial cues, object dynamics, and scene changes—will be key to achieving more grounded and effective visual decision-making. In terms of model architecture, improvements often stem from better attention to visual details that align with long-horizon reasoning steps. Spatial-temporal modeling offers unique opportunities for more creative action definition to manipulate the visual information for better reasoning. Similarly, during data curation, verifiers commonly depend on grounding models or language cues—indicating potential for more creative, vision-focused verifier designs.
@@ 3.1 Policy Optimization
To solve the above reward maximization problem, recent advancements in visual reasoning have prominently focused on integrating reinforcement learning (RL) and imitation learning (IL) to align MLLMs more closely with human reasoning. IL approaches, such as Thought Cloning (2503.08525-wei_gtr:_2025), surpass traditional cloning by aligning intermediate reasoning steps instead of final outputs alone. Coupled with iterative methods like DAgger, this reduces dataset shift and hallucination. LLaVA-Critic (2410.02712-xiong_llava-critic:_2024) further generalizes reward signals via preference-based alignment, boosting multimodal reasoning effectiveness with minimal modality-specific fine-tuning. RL from Simulations (RLS3) (2501.18880-waite_rls3:_2025), employing Soft Actor-Critic (SAC) agents, significantly enhances spatial reasoning accuracy and efficiency in models like PaliGemma, outperforming random exploration baselines (2501.18880-waite_rls3:_2025). FuRL (2406.00645-fu_furl:_2024) addresses RL reward misalignment through dual-stage alignment, effectively mitigating sparse reward challenges. Adaptive Reward Programming (ARP-DT+) (2309.10790-kim_guide_2023), integrating CLIP representations with Decision Transformers, enhances reasoning robustness across domains by employing active inference to counter distractions and domain shifts.
@@ 3.2 Reward Alignment
Recent approaches have explored iterative refinement and reflection mechanisms, token-level rewards, and automated benchmarking to enhance reasoning quality and mitigate hallucinations. PARM++ (2501.13926-guo_can_2025) introduces a reflection-based mechanism that iteratively refines outputs by identifying misalignments and actively correcting them through self-assessment loops, significantly improving test-time GenEval scores up to three reflection cycles (2501.13926-guo_can_2025). Similarly, REVERIE employs rationale-based training, explicitly supervising reasoning steps, thus improving coherence and reducing hallucinations compared to models without rationales (2407.11422-zhang_reflective_2024). Fine-grained reward models, such as MMViG (2402.06118-yan_vigor:_2024) and Reward Alignment via Preference Learning (RAPL) (2310.07932-tian_what_2024), emphasize granular feedback at individual reasoning steps, achieving precise improvements without manual reward tuning. Similarly, CLIP-DPO (2408.10433-ouali_clip-dpo:_2024) leverages pre-trained CLIP directly as a reward function, streamlining reward modeling. Challenges such as constrained perceptual fields are addressed through dynamic evaluation strategies like zooming and shifting, emphasizing the importance of sequential decision-making in visual reasoning (2410.04659-wang_actiview:_2024). Structural issues like positional bias and patch-boundary effects further underscore the need for targeted architectural optimizations (2402.07384-zhang_exploring_2024). Recent approaches integrate reward signals directly into generation. TLDR (2410.04734-fu_tldr:_2024) applies token-level binary rewards, enabling real-time feedback to mitigate hallucinations and boost both annotation efficiency and backbone model performance (2410.04734-fu_tldr:_2024). Complementarily, EACO enhances intermediate object recognition, significantly reducing hallucinations compared to earlier models like LLaVA-v1.6-7B (2412.04903-wang_eaco:_2024). Distillation frameworks also contribute: FIRE leverages iterative student–teacher feedback and structured evaluation to refine visual reasoning (2407.11522-li_fire:_2024), while SILKIE demonstrates the scalability of GPT-4V-generated feedback in improving perception and cognition without human intervention (2312.10665-li_silkie:_2023).
@@ 3.3 Model Architecture
Nested architectures, such as MaGNeTS, have emerged as effective means of balancing computational efficiency with model accuracy by employing parameter-sharing and caching mechanisms during inference (2502.00382-goyal_masked_2025). Intrinsic activation approaches, notably ROSS, bypass external modules by integrating multimodal understanding directly into the model’s core, yielding improved adaptability to diverse visual inputs, including depth maps (2410.09575-wang_reconstructive_2024). Additionally, DIFFLMM incorporates diffusion models and attention-based segmentation, enhancing visual grounding precision (2410.08209-cao_emerging_2024), while Mini-Monkey targets the ”sawtooth effect” in lightweight models through adaptive cropping and scale compression, achieving superior visual-text alignment without extensive computational resources (2408.02034-huang_mini-monkey:_2024). SEA dynamically adjusts embedding alignment strategies to maintain performance across varying resolutions and model sizes (2408.11813-yin_sea:_2024). Concurrently, PAE-LaMM demonstrates the combined effectiveness of vision encoder adaptation and pixel reconstruction tasks, systematically enhancing visual detail awareness and question-answering performance (2408.03940-gou_how_2024). Conversely, LLAVIDAL and EventLens uniquely integrate domain-specific visual elements, such as object-centric embeddings and event-aware tokens, respectively, directly within LLM architectures to effectively reason about activities and temporal contexts (2406.09390-reilly_llavidal:_2024; 2404.13847-ma_eventlens:_2024).
@@ 3.4 Spatial Temporal Modeling
Unlike pure language reasoning, visual reasoning involves a richer action space, allowing operations such as zooming, cropping, and frame selection, which enables the model to interact with and manipulate visual inputs. The GeoGLIP pipeline utilizes geometric pre-training with a dynamic feature router to advance fine-grained visual mathematical understanding by balancing symbolic and visual modalities based on task demands (2501.06430-zhang_open_2025). Concurrently, hallucination issues are mitigated via targeted fine-tuning and prompting strategies grounded in recent vision-language work like CapQA (2501.02964-hu_socratic_2025). For spatial reasoning, simpler graph topologies have been shown to enhance performance, while structural complexity reduces accuracy in edge-centric tasks (2412.13540-zhu_benchmarking_2024). Augmentation via uncertainty-aware active inference further boosts attribute detection precision (2412.07012-zhang_provision:_2024). Intrinsic spatial-temporal modeling is advanced through architectures embedding multi-modal understanding without reliance on external depth tools (2410.09575-wang_reconstructive_2024), and dual-branch structures improve video temporal grounding in multi-hop reasoning (2408.14469-chen_grounded_2024). The shift toward multimodal benchmarks better reflects real-world spatial-temporal reasoning (2408.08632-li_survey_2024). Integrating rationalization with answer generation enhances temporal reasoning (2407.06189-zohar_video-star:_2024), while automated annotation improves data scalability and bridges gaps between commercial and open-source Video-LLMs (2406.11303-li_videovista:_2024). Multi-modal integration and active inference optimize modality selection for multi-step spatial-temporal reasoning (2405.03272-zhang_worldqa:_2024). Adjusting learning objectives to minimize hallucinations improves precision and EOS decision-making (2402.14545-yue_less_2024).
@@ 4 Test-Time compute
ȷ Search Strategies Current search algorithms primarily operate on distinct textual tokens. However, innovative approaches are emerging for raw visual tokens, termed Creative Visual Search, expanding search capabilities directly within visual spaces. Moreover, Multi-Granularity Spatial Search is gaining attention as an effective extension, enhancing exploration across diverse spatial granularities. Reward and Feedback Reward systems can significantly benefit from deeper visual insights. Leveraging visual prompts encourages models to engage in self-reflection grounded in visual information, substantially improving semantic alignment and decision-making effectiveness. Iterative Refinement Determining the alignment between textual tokens and visual information remains challenging. Consequently, model self-reflection can produce hallucinated interpretations. Nevertheless, these reflections offer valuable insights, highlighting new pathways to mitigate unnecessary or hallucinated reflections and enhance model reliability.
@@ 4.1 Search Strategies
Monte Carlo Tree Search (MCTS) has emerged as a powerful tool for managing uncertainty and sequential decision-making. 2501.10674-imam_can_2025 applied MCTS to enhance temporal reasoning, reducing hallucinations in time-dependent data. Building on this, CoMCTS(2412.18319-yao_mulberry:_2024) unified multiple MLLMs within a collaborative tree search, addressing model bias and improving accuracy and efficiency. Vision-specific strategies advanced with VisVM (2412.03704-wang_scaling_2024), which estimates long-term candidate value to reduce hallucinations, outperforming immediate-alignment methods like CLIP-PRM. Video reasoning similarly benefited from reward modeling and adaptive exploration to prioritize key frames (2412.10471-yang_vca:_2024). Structured search methods also gained traction. LLaVA-o1 (2411.10440-xu_llava-cot:_2025) uses stage-level beam search for complex reasoning, while MVP (2408.17150-qu_look_2024) aggregates certainty across multi-perspective captions to resist adversarial inputs. DC2 (2408.15556-wang_divide_2024) applies MCTS-based cropping to focus on salient image regions for high-resolution reasoning. Multimodal and temporal search frameworks like VideoVista (2406.11303-li_videovista:_2024), WorldRetriver (2405.03272-zhang_worldqa:_2024), and DynRefer (2405.16071-zhao_dynrefer:_2024) surpass static baselines using adaptive sampling, fusion, and stochastic inference. Step-by-step comparative reasoning with LMs also enhances video QA (2404.16222-nagarajan_step_2024). Innovations in context refinement (VURF (2403.14743-mahmood_vurf:_2024)), image decomposition (V*(2312.14135-wu_v*:_2023)), and dynamic tool use (AVIS(2306.08129-hu_avis:_2023)) highlight the shift toward adaptive visual reasoning. FuRL (2406.00645-fu_furl:_2024) aligns reward modeling with iterative fine-tuning for better performance. In embodied AI, combining chain-of-thought, self-verification, and MCTS-driven planning enables scalable, robust decision-making in dynamic environments (2404.15190-shin_socratic_2024).
@@ 4.2 Adaptive Inference
Adaptive inference is reshaping vision-language reasoning by enabling dynamic, context-sensitive processing that improves both accuracy and efficiency. Central to this is the iterative evaluation and refinement of outputs, often through internal feedback and external verification. LOOKBACK enhances correction accuracy through iterative visual re-examination within each reasoning step (2412.02172-wu_visco:_2024). IXC-2.5 balances performance and response length via Best-of-N sampling guided by reward models (2501.12368-zang_internlm-xcomposer2.5-reward:_2025), while PARM++ uses reflection-based refinement to align generated images with prompts (2501.13926-guo_can_2025). Further innovations include ProgressCaptioner’s sliding-window approach for tracking action progress over time (2412.02071-xue_progress-aware_2024). To combat multi-modal hallucinations, MEMVR triggers visual retracing based on uncertainty, optimizing cost-accuracy tradeoffs (2410.03577-zou_look_2024), while MVP aggregates certainty across diverse views (2408.17150-qu_look_2024). SID applies token-level contrastive decoding to filter irrelevant content (2408.02032-huo_self-introspective_2024). Models like DualFocus integrate macro and micro reasoning for fine-grained attention control (2402.14767-cao_dualfocus:_2024), and LISA++ enables test-time compute scaling without retraining (2312.17240-yang_lisa++:_2024). PerceptionGPT accelerates inference via adaptive token embeddings that encode spatial cues and dynamically weigh layers (2311.06612-pi_perceptiongpt:_2023).
@@ 4.3 Reward
2501.13926-guo_can_2025 introduce ORM and PRM, showing that ORM’s final-step evaluation significantly enhances image generation through ranking data. PARM++ further refines this method by enabling iterative reflection for self-correction, thus improving prompt fidelity. Complementarily, 2501.05444-hao_can_2025 highlight persistent errors in VLM spatial reasoning and explanation, proposing reward re-ranking as an interim solution. To assess VL-GenRMs more comprehensively, 2411.17451-li_vlrewardbench:_2024 propose VL-RewardBench, pinpointing perception and reasoning challenges while advocating critic training and co-evolutionary reward learning. Self-training methods also gain prominence: R3V (2411.00855-cheng_vision-language_2024) leverages synthesized rationales to enhance noise-robust Chain-of-Thought (CoT) refinement, while CECE (2410.22315-cascante-bonilla_natural_2024) enriches evaluation using LLM-generated entailment/contradiction captions. Enhanced supervision methods also emerge. 2410.16198-zhang_improve_2024 enhance CoT reasoning through ShareGPT-4O distillation, Supervised Fine-Tuning (SFT), and Direct Preference Optimization (DPO). LLaVA-Critic (2410.02712-xiong_llava-critic:_2024) builds upon this using critic-guided DPO without external human feedback. FuRL (2406.00645-fu_furl:_2024) introduces a two-stage tuning approach addressing misalignments and temporal errors, while HYDRA (2403.12884-ke_hydra:_2024) employs dynamic loops for adaptive instruction ranking. Additionally, m&m’s (2403.11085-ma_m&ms:_2024) provides a structured framework for evaluating multi-step planning via tool and argument accuracy.